# CNSocialDepress: A Chinese Social Media Dataset for Depression Risk Detection and Structured Analysis

### v1

@@ 4.1 Module I: Dimension-Wise Automatic Labeling
    Texts from positive users in CNSD Gold containing depression-related semantic segments for each dimension $D_{k}$ were extracted, forming sets $S_{k}$.
@@
    The dimension with the fewest samples ($n_{\text{min}}=442$, corresponding to "depression-related medical expressions") was identified:
    $n_{\text{min}}=\min_{k\in\{1,2,\dots,6\}}n_{k}=442$
@@
    To balance data across dimensions, texts for each dimension $D_{k}$ were downsampled to $n_{\text{min}}$ if $n_{k}>n_{\text{min}}$, forming balanced sets $S^{\prime}_{k}$:
    $S^{\prime}_{k}=\begin{cases}S_{k},&\text{if }n_{k}=n_{\text{min}},\\ \text{RandomSubset}(S_{k},n_{\text{min}}),&\text{if }n_{k}>n_{\text{min}}.\end{cases}$
    This resulted in $6\times 442$ labeled positive texts. Each text is annotated in the format: “This text segment belongs to [category] because it references [feature],” supplemented by 20 randomly generated expressions sharing the same meaning (see Appendix A.2).
@@
    For negative data, posts from negative users in CNSD Gold were split, and texts with any depression-related dimension were filtered out. From the remaining unlabeled texts, $6\times 442$ were uniformly sampled and assigned one of 20 distinct label expressions indicating absence of negative emotions (details in Appendix A.3).
@@
    The Module I model (Qwen2.5-14B) was then fine-tuned using LoRA Hu et al. (2022) on this combined set of $2\times(6\times 442)$ texts for 1 epoch at a learning rate of 5e-05 (details in Appendix A.4).
4.2 Module II: Automatic Verification and Summarization
In the second module, we applied the fine-tuned model from Module I to label every text for each user, extracting all texts that contain any depression-related semantics. We then leveraged the Deepseek R1 671B model DeepSeek-AI et al. (2025) to verify the correctness of both labels and texts on a case-by-case basis. Following this verification step, we employed a few-shot approach to summarize the extracted texts. In the few-shot setting, we instruct the model to assume the role of experts in psychology and text sentiment analysis. Additionally, we provide the model with detailed task requirements and exemplary gold-standard answer samples. Details are in the Appendix 3.
5 Experiments
We design our experiments around two central research questions. First , how effectively our data generation pipeline in producing high-quality structured annotations? Second , how well do various large language models (LLMs) perform on tasks involving user-level depression risk summarization and structured psychological analysis?
Based on these questions, our experiments aim to:
@@ 5 Experiments
    To assess the quality of text generated for six-dimensional depression analysis by models fine-tuned on CNSD Gold and CNSD Silver, comparing them against baseline models and few-shot approaches.
@@
    To evaluate the performance of these fine-tuned models on the downstream task of user-level depression risk classification.
@@
    To benchmark the performance of various leading LLMs on a structured depression analysis summarization task using our CNSD Gold dataset.
Additionally, as shown in the Appendix 5.5, we demonstrate the use of the CNSD Gold dataset as the test set for the binary depression classification task, further illustrating the dataset’s versatility.
5.1 Baseline Models
In the following experiments, we employ DeepSeek-R1 Distill-14B DeepSeek-AI et al. (2025), Qwen2.5-14B QwenTeam (2024), GPT-4o  [^5] https://openai.com/index/hello-GPT-4o/, GPT-4o-Mini [^6] https://openai.com/index/GPT-4o-mini-advancing-cost-efficient-intelligence/, DeepSeek-R1-671B [^7] https://www.deepseek.com/ DeepSeek-AI et al. (2025), and Llama3-8B-Chinese-Chat Wang et al. (2024b) as our generative baseline models. We summarize the key characteristics of each model on A.1.
5.2 Experimental Setup
For our experiments, we used NVIDIA A800 80G and A100 80G GPUs. Unless otherwise specified, all generation experiments were conducted with a temperature setting of 0.7. All fine-tuning experiments were conducted using the LLama Factory framework Zheng et al. (2024), employing the LoRA method Hu et al. (2022) for one epoch with a learning rate of 5e-5.
Our experiments consist of two main parts. The first part evaluates the performance of our pipeline on data generation, while the second part presents benchmarks on CNSD Gold dataset for Structured Summarization and Analysis Generation.
In our classification tasks, we use accuracy, recall, precision, and F1-score as evaluation metrics. For text generation experiments, we employ BLEU Papineni et al. (2002), ROUGE-1 Lin (2004), and BERTScoreZhang et al. (2020) for performance assessment.
5.3 Task I: Data Generation
In this section, we conducted multiple experiments on the CNSD Test dataset. We primarily evaluated the Qwen2.5 14B models fine-tuned on CNSD Gold and CNSD Silver:
  • Qwen2.5 14B Gold: Fine-tuned using the CNSD Gold.
  • Qwen2.5 14B Silver: Fine-tuned using the CNSD Silver. which was generated by applying the pipeline described in Section 4 to 100 randomly selected positive users and 100 randomly selected negative users from the SWDD binary classification dataset.
We compared the performance of these models with other baselines in two aspects:
@@ 5.3 Task I: Data Generation
    The performance of the depression risk classification.
@@
    The quality of the generated text in the six-dimensional analyses section (including text hallucinations).
Throughout these experiments, unless noted otherwise, we employed a user-level generation strategy. This approach involves providing the complete set of a user’s Weibo posts along with an instruction that requires the model to directly output the user’s depression label along with the corresponding justification.
@@
Table 3: Comparison of Generation Strategies on Text Quality: In the table, FS stands for Few-Shot, and Pipeline is based on the Automated Dataset Generation Pipeline we proposed in Section 4 and our proposed dataset in Section 3. The prompt used in this experiment can be found in Appendix A.6. The purpose is to demonstrate that, using the same prompt, our pipeline achieves the highest text generation quality that meets our task requirements.
5.3.1 Six-Dimensional Analysis of Depression task
Research has demonstrated that structured knowledge enhances the benefits of integrating structured data into LLMs Moiseev et al. (2022), thereby improving output quality. CNSD Gold and CNSD Silver are datasets with fine-grained structured annotations across six dimensions to summarize and analyze social media users. Fine-tuning on these datasets should reduces textual hallucinations and achieves text generation quality comparable to the other much larger model.
@@ 5.3.1 Six-Dimensional Analysis of Depression task
Table 3 shows the comparisons between different generation strategies in terms of the quality of the produced text. Notably, our Pipeline achieves the highest scores across all three metrics:BERTScore (0.791), ROUGE-1 (0.478), and BLEU (0.288), indicating superior semantic alignment, lexical overlap, and n-gram precision, respectively. In contrast, with the FS (few-shot) strategies, all models: Qwen2.5 14B, GPT-4o, and DeepSeek R1 671B, yield comparatively lower performance on each metric. Even the large-scale DeepSeek R1 671B model does not surpass the Pipeline approach, suggesting that the proposed Pipeline method effectively leverages its generation procedure to produce higher-quality text that more closely matches the reference, outperforming both smaller and larger models under the few-shot setting.
Additionally, we also report the comparison of generation quality between the original model, Qwen2.5 14B Gold, and Qwen2.5 14B Silver. (Table 5)
For automatic metrics, the original model achieved a Bert-Score of 0.649, which increased to 0.764 for Gold (+16.9%) and 0.787 for Silver (+21.5%), showing stronger semantic alignment. Rouge-1 rose from 0.076 to 0.217 for Gold (+172%) and 0.218 for Silver (+172%), indicating improved lexical overlap. Bleu increased from 0.075 to 0.186 for Gold (+149%) and 0.237 for Silver (+218%), reflecting better n-gram precision and structural quality.
The human evaluation was conducted by two linguistic experts across accuracy, coverage, and hallucination. Compared to the original Qwen2.5 14B, both Silver and Gold showed clear improvements. Silver improved by about 33.4% in accuracy, 37.1% in coverage, and 15.5% in hallucination. Gold achieved even stronger gains, with 42.9% in accuracy, 58.0% in coverage, and 26.8% in hallucination. Overall, Gold performed best in human evaluation, while Silver, though slightly behind, still showed significant gains over the original and reached levels close to Gold.
@@
In summary, Gold achieved the strongest results in human evaluation, while Silver performed better in automatic metrics. Both models clearly outperformed the original, showing that fine-tuning with our dataset or pipeline can significantly enhance generation quality at semantic, lexical, and structural levels.
Model/Test Dataset Swdd Origin Wu3d Ours (Gold)
Acc F1 Acc F1 Acc F1
RoBERTa Chinese (trained by swdd origin) 0.89 0.90 0.76 0.80 0.88 0.87
RoBERTa Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.82 0.78
Bert based Chinese (trained by wu3d) 0.88 0.87 0.90 0.91 0.83 0.80
StructBERT-mental (trained by wu3d) 0.85 0.85 0.92 0.92 0.83 0.80
Llama3-8b 0.86 0.86 0.79 0.79 0.60 0.63
Glm4-9b-chat 0.76 0.81 0.77 0.81 0.79 0.82
Qwen2.5 7b 0.88 0.89 0.84 0.85 0.94 0.94
GPT-4o-Mini 0.65 0.74 0.65 0.74 0.53 0.68
GPT-4o 0.92 0.92 0.91 0.92 0.83 0.86
DeepSeek-R1 671b 0.86 0.87 0.89 0.90 0.81 0.84
Model Bert-Score Rouge1 Bleu H.acc H.cov H.hallu
Qwen2.5 14B 0.649 0.076 0.075 0.583 0.574 0.657
Qwen2.5 14B Silver 0.787 0.218 0.237 0.778 0.787 0.759
Qwen2.5 14B Gold 0.764 0.217 0.186 0.833 0.907 0.833
@@
Table 5: Comparison of text quality across different models. ’Qwen2.5 14B Gold’ denotes the model fine-tuned on our proposed dataset, while ’Qwen2.5 14B Silver’ denotes the model fine-tuned using data generated from the pipeline strategy described in Table 3, with further details in Section 4. The table reports three human evaluation metrics: acc (accuracy of the generated text), cov (content coverage), and hallu (hallucination). Higher values across all metrics indicate better generation performance.
5.3.2 Classification task
@@
In the depression classification task, Qwen2.5 14B Silver achieves the best performance compared to other models, with an accuracy of 0.944 and an F1 score of 0.941, even surpassing Qwen2.5 14B Gold, which was fine-tuned on human-annotated data. Furthermore, it outperforms large-scale models such as GPT-4o (accuracy = 0.917, F1 score = 0.923) and DeepSeek R1 671B (accuracy = 0.861, F1 score = 0.872). This result provides evidence for the effectiveness of our generation pipeline and the high quality of the data produced through this pipeline. The details of the experimental results are in the Table 6.
@@
Qwen2.5 14B 0.889 0.944 0.850 0.894
Qwen2.5 14B_FT_Silver 0.944 0.889 1.000 0.941
Qwen2.5 14B_FT_Gold 0.861 0.889 0.842 0.864
@@
GPT-4o Mini 0.667 1.000 0.600 0.750
DeepSeek R1 671B 0.861 0.944 0.801 0.872
DeepSeek R1 Distill-14B 0.889 1.000 0.818 0.900
@@
5.4 Task II: Structured Summarization and Analysis Generation
The proposed dataset effectively supports research on generative models in the field of internet depression risk analyses, demonstrating particular value for user depressive mental state summarization tasks. Due to the absence of comparable open-source datasets, we conducted preliminary evaluations on mainstream generative models using our annotated dataset containing 233 user-level samples. The detailed results are demonstrated in the table 7. We used the basic prompt as follows:
Please carefully read the following multiple texts and then answer the question: Based on these texts, does this user exhibit a depressive mood? If the user exhibits depressive mood, answer ’Yes’; if the user does not exhibit depressive mood, answer ’No’. Then, provide your step-by-step analysis process.
Model Bert-Score Rouge1 Bleu
@@ 5.4 Task II: Structured Summarization and Analysis Generation
Table 7: Performance Comparison on Depression Risk Analysis Summarization Generation Task.
The results reveal variations across models in the task of risk analysis summarization generation:
Bert-Score: Narrow range (0.65-0.71) with GPT-4o leading at 0.71, indicating superior semantic understanding. Rouge1: DeepSeek-R1-671B achieves remarkable 0.45 score, suggesting 10% improvement over second-best performer. Bleu: Maximum 15.17 from GPT-4o vs minimum 9.36 from Qwen2.5-7b, reflecting lexical alignment differences.
5.5 Task III: Classification Experimentations
In the depression risk binary classification task, we used three datasets, including the original SWDD dataset Cai et al. (2023) and the WU3D dataset Wang et al. (2020). Both of these datasets were randomly sampled with 200 negative users and 200 positive users for the experiment. Details provided in the table 4.
In the classification experiments, in addition to the generative models mentioned in the section A.1, we also used models based on the BERT architecture with the details on A.7.
6 Conclusion

### v2

@@ 4.1.  Module I: Dimension-Wise Automatic Labeling
    For each dimension $D_{k}$, we extract posts from positive users in CNSD Gold that contain at least one annotated span of $D_{k}$. These posts form a set $S_{k}$.
@@
    We identify the dimension with the fewest training posts. The minimum count is $n_{\text{min}}=442$, which corresponds to depression-related medical expressions:
    $n_{\text{min}}=\min_{k\in\{1,2,\dots,6\}}n_{k}=442.$
@@
    To balance dimensions, we downsample each $S_{k}$ to $n_{\text{min}}$ when $n_{k}>n_{\text{min}}$. This yields balanced sets $S^{\prime}_{k}$:
    $\small S^{\prime}_{k}=\begin{cases}S_{k},&\text{if }n_{k}=n_{\text{min}},\\ \text{RandomSubset}(S_{k},n_{\text{min}}),&\text{if }n_{k}>n_{\text{min}}.\end{cases}$
    We obtain $6\times 442$ labeled positive posts. Each post is formatted using an instruction template such as: “This post belongs to [category] because it mentions [evidence].” We further augment the labels with 20 paraphrased expressions with the same meaning (Appendix A.2).
@@
    For negative training data, we split posts from negative users in CNSD Gold into individual texts. We remove any text that contains depression-related content in any dimension. From the remaining texts, we uniformly sample $6\times 442$ examples. We assign them negative label expressions indicating that there is no evidence for the target dimension (As with the positive examples, we create 20 negative label expressions to reduce overfitting; details are provided in A.3).
@@
    We fine-tune Qwen2.5-14B with LoRA Hu et al. (2022) on the combined set of $2\times(6\times 442)$ texts for 1 epoch. We use a learning rate of $5\times 10^{-5}$, with the prompt provided in Appendix A.4.
4.2.  Module II: Automatic Verification and Summarization
In Module II, we apply the fine-tuned Module I model to label all posts for each user. We then collect posts that are assigned to at least one dimension. Next, we use DeepSeek-R1-671B DeepSeek-AI et al. (2025) to verify the assigned labels and the supporting evidence on a case-by-case basis. After verification, we summarize the collected posts using a few-shot prompting setup. The prompt instructs the model to act as a psychology expert and to produce structured summaries aligned with CNSD Gold. We also provide detailed requirements and gold-standard examples. The full prompt is shown in Appendix A.7.
5.  Experiments
We design our experiments to answer two research questions. First, how effectively does our data generation pipeline produce high-quality structured annotations ? Second, how well do various large language models (LLMs) perform on user-level depression risk summarization and structured psychological analysis ?
Based on these questions, we conduct three sets of experiments:
@@ 5.  Experiments
    To assess the quality of model-generated six-dimensional structured depression analyses. We compare models fine-tuned on CNSD Gold and CNSD Silver with baseline models and few-shot prompting.
@@
    To evaluate the performance of these fine-tuned models on user-level depression risk classification.
@@
    To benchmark leading LLMs on a structured depression analysis summarization task using CNSD Gold.
In Section 5.5, we also use CNSD Gold as a test set for binary depression classification. This further illustrates the dataset’s versatility.
5.1.  Baseline Models
We use the following generative baselines: DeepSeek-R1-Distill-14B DeepSeek-AI et al. (2025), Qwen2.5-14B QwenTeam (2024), GPT-4o [^4] https://openai.com/index/hello-GPT-4o/, GPT-4o-mini [^5] https://openai.com/index/GPT-4o-mini-advancing-cost-efficient-intelligence/, DeepSeek-R1-671B [^6] https://www.deepseek.com/ DeepSeek-AI et al. (2025), and Llama3-8B-Chinese-Chat Wang et al. (2024b). We summarize key characteristics in Table A.1.
5.2.  Experimental Setup
We use NVIDIA A800 (80GB) and A100 (80GB) GPUs. Unless otherwise specified, we set the generation temperature to 0.7. All fine-tuning experiments use the LLaMA-Factory framework Zheng et al. (2024). We apply LoRA Hu et al. (2022) for 1 epoch with a learning rate of $5\times 10^{-5}$.
Our experiments have two parts. The first evaluates the pipeline for data generation. The second benchmarks models on the CNSD Gold dataset for structured summarization and analysis generation.
For classification, we report accuracy, precision, recall, and F1-score. For generation, we report BLEU Papineni et al. (2002), ROUGE-1 Lin (2004), and BERTScore Zhang et al. (2020).
5.3.  Task I: Data Generation
We conduct experiments on the CNSD Test dataset. We focus on two Qwen2.5-14B models fine-tuned on CNSD Gold and CNSD Silver:
  • Qwen2.5-14B Gold: Fine-tuned on CNSD Gold.
  • Qwen2.5-14B Silver: Fine-tuned on CNSD Silver. CNSD Silver is generated using the pipeline in Section 4 on 100 positive and 100 negative users sampled from SWDD.
We compare these models with other baselines in two aspects:
@@ 5.3.  Task I: Data Generation
    Depression risk classification performance.
@@
    Quality of the generated six-dimensional analyses, including hallucinated or unsupported content.
Unless noted otherwise, we use user-level generation. We concatenate all posts from a user as input. We instruct the model to output a binary label and a corresponding justification.
@@
Table 3: Comparison of generation strategies on text quality. In the table, FS stands for Few-Shot. Pipeline refers to the automated dataset generation pipeline proposed in Section 4, applied to our proposed dataset in Section 3. The purpose is to show that, with the same prompt, our pipeline achieves the highest text generation quality while meeting our task requirements. The prompt used in this experiment is shown in Figure 3.
Figure 3: Prompt used for Table 3.
5.3.1.  Six-Dimensional Structured Depression Analysis
Prior work suggests that incorporating structured knowledge can improve LLM outputs Moiseev et al. (2022). CNSD Gold and CNSD Silver provide fine-grained structured annotations across six dimensions. They support user-level summarization and analysis of social media content. We hypothesize that fine-tuning on these datasets can reduce hallucinations. It may also improve generation quality compared to few-shot prompting and larger models.
@@ 5.3.1.  Six-Dimensional Structured Depression Analysis
Table 3 compares different generation strategies using automatic metrics. Our Pipeline achieves the best scores on all three metrics: BERTScore (0.791), ROUGE-1 (0.478), and BLEU (0.288). These results indicate stronger semantic alignment, higher lexical overlap, and better $n$-gram precision. Few-shot (FS) prompting yields lower scores across models, including Qwen2.5-14B, GPT-4o, and DeepSeek-R1-671B. Notably, DeepSeek-R1-671B does not outperform the Pipeline. This suggests that the proposed Pipeline produces text that matches the reference more closely than few-shot prompting, across both smaller and larger models.
We also compare the original Qwen2.5-14B model with Qwen2.5-14B Gold and Qwen2.5-14B Silver (Table 4). We use two complementary evaluation methods.
  • Automatic metrics. The original model achieves a BERTScore of 0.649. It increases to 0.764 for Gold (+16.9%) and 0.787 for Silver (+21.5%). ROUGE-1 rises from 0.076 to 0.217 for Gold (+172%) and 0.218 for Silver (+172%). BLEU improves from 0.075 to 0.186 for Gold (+149%) and 0.237 for Silver (+218%).
  • Human evaluation. Two linguistic experts evaluate accuracy, coverage, and hallucination. Compared to the original Qwen2.5-14B, both Gold and Silver improve substantially. Silver increases by 33.4% in accuracy and 37.1% in coverage. Its hallucination score increases by 15.5% (higher is better). Gold achieves larger gains. Accuracy increases by 42.9% and coverage increases by 58.0%. The hallucination score increases by 26.8%. Overall, Gold performs best in human evaluation. Silver is slightly behind but remains close to Gold and clearly outperforms the original model.
@@
In summary, Gold performs best in human evaluation. Silver performs better on automatic metrics. Both models substantially outperform the original model. These results show that fine-tuning with our dataset and pipeline can improve generation quality at semantic, lexical, and structural levels.
Model BERTScore ROUGE-1 BLEU Human.acc Human.cov Human.hallu
Qwen2.5-14B 0.649 0.076 0.075 0.583 0.574 0.657
Qwen2.5-14B Silver 0.787 0.218 0.237 0.778 0.787 0.759
Qwen2.5-14B Gold 0.764 0.217 0.186 0.833 0.907 0.833
@@
Table 4: Comparison of generation quality across models. Qwen2.5-14B Gold is fine-tuned on CNSD Gold. Qwen2.5-14B Silver is fine-tuned on CNSD Silver, which is generated by our pipeline (Section 4). We report automatic metrics (BERTScore, ROUGE-1, BLEU) and human evaluation metrics: acc (accuracy), cov (content coverage), and hallu (hallucination; higher is better). Higher values indicate better performance for all metrics.
5.3.2.  Classification Task
@@
In the depression classification task, Qwen2.5-14B Silver achieves the best performance among all models. It reaches an accuracy of 0.944 and an F1 score of 0.941. It also surpasses Qwen2.5-14B Gold, which is fine-tuned on human-annotated data. Moreover, it outperforms large-scale models such as GPT-4o (accuracy = 0.917, F1 = 0.923) and DeepSeek-R1-671B (accuracy = 0.861, F1 = 0.872). These results support the effectiveness of our generation pipeline and suggest that the resulting silver data are of high quality. Detailed results are reported in Table 5.
@@
Qwen2.5-14B 0.889 0.944 0.850 0.894
Qwen2.5-14B (FT, Silver) 0.944 0.889 1.000 0.941
Qwen2.5-14B (FT, Gold) 0.861 0.889 0.842 0.864
@@
GPT-4o-mini 0.667 1.000 0.600 0.750
DeepSeek-R1-671B 0.861 0.944 0.801 0.872
DeepSeek-R1-Distill-14B 0.889 1.000 0.818 0.900
@@
5.4.  Task II: Structured Summarization and Analysis Generation
Our dataset supports research on generative models for depression risk assessment from social media. It is particularly useful for user-level summarization of depressive signals. To the best of our knowledge, there is no directly comparable open-source dataset with the same annotation schema. We therefore conduct an initial benchmark of mainstream generative models on our annotated dataset of 233 users. Results are reported in Table 6.
We use the following prompt:
  • Instruction: Read the following texts written by one user. Then answer the question.
  • Question: Based on these texts, does this user exhibit a depressive mood ?
    • If the user exhibits a depressive mood, answer Yes.
    • If the user does not exhibit a depressive mood, answer No.
  • Output requirements: Provide a brief justification based on evidence from the texts. Then produce a structured summary aligned with the annotation schema.
The results vary across models. BERTScore falls in a narrow range (0.654–0.710), with GPT-4o achieving the highest score. ROUGE-1 is highest for DeepSeek-R1-671B (0.454), which exceeds the second-best model (GPT-4o at 0.346) by an absolute margin of 0.108. BLEU ranges from 0.093 (Qwen2.5-7B) to 0.152 (GPT-4o).
Model BERT Score ROUGE-1 BLEU
@@ 5.4.  Task II: Structured Summarization and Analysis Generation
Table 6: Model performance on the depression risk summarization and analysis generation task.
$5.$  Task III: Classification Experiments
Our dataset is primarily designed for generation tasks. We also conduct a binary classification experiment for depression risk. In addition to our dataset, we use the original SWDD and WU3D datasets. For each dataset, we randomly sample 200 positive and 200 negative users.
Besides the generative models described in Appendix A.1, we evaluate BERT-based classifiers (Devlin et al., 2019). Implementation details are provided in Appendix A.5. Results are reported in Appendix A.6.
6.  Conclusion